# GlobalGeoTree: A Multi-Granular Vision-Language Dataset for Global Tree Species Classification

### v2

@@ Abstract
Global tree species mapping using remote sensing data is vital for biodiversity monitoring, forest management, and ecological research. However, progress in this field has been constrained by the scarcity of large-scale, labeled datasets. To address this, we introduce GlobalGeoTree—a comprehensive global dataset for tree species classification. GlobalGeoTree comprises 6.3 million geolocated tree occurrences, spanning 275 families, 2,734 genera, and 21,001 species across the hierarchical taxonomic levels. Each sample is paired with Sentinel-2 image time series and 27 auxiliary environmental variables, encompassing bioclimatic, geographic, and soil data. The dataset is partitioned into GlobalGeoTree-6M for model pretraining and curated evaluation subsets, primarily GlobalGeoTree-10kEval for zero-shot and few-shot benchmarking. To demonstrate the utility of the dataset, we introduce a baseline model, GeoTreeCLIP, which leverages paired remote sensing data and taxonomic text labels within a vision-language framework pretrained on GlobalGeoTree-6M. Experimental results show that GeoTreeCLIP achieves substantial improvements in zero- and few-shot classification on GlobalGeoTree-10kEval over existing advanced models. By making the dataset, models, and code publicly available, we aim to establish a benchmark to advance tree species classification and foster innovation in biodiversity research and ecological applications.
@@
1 Introduction
Forests cover approximately 31% of the global land surface [1] and provide essential ecosystem services, including carbon sequestration [2], biodiversity conservation [3], and climate regulation [4]. Accurate and large-scale mapping of tree species plays an increasingly vital role in addressing pressing environmental challenges [5], including effective biodiversity monitoring [6], informed forest management practices [7], and comprehensive ecological research aimed at understanding the complex impacts of climate change [8].
Traditional ground-based forest monitoring methods [9], while providing detailed information, are often limited in their spatial and temporal coverage, making it challenging to obtain a comprehensive understanding of global forest composition and dynamics. In contrast, remote sensing has emerged as a key technology for large-scale forest monitoring, offering non-invasive and cost-effective approaches to tree species classification [10]. Despite significant advancements in this field, progress has been constrained by the limited availability of comprehensive, high-quality, and accurately labeled datasets that capture the global diversity of tree species [11]. Existing datasets typically focus on specific geographic regions or limited taxonomic coverage, hampering the development of models with global applicability [12].
@@ 1 Introduction
In addition to the dataset, we introduce GeoTreeCLIP, a vision-language model specifically designed for tree species classification. Drawing on frameworks like CLIP [13], our approach aligns satellite imagery with taxonomic labels to learn nuanced representations. Unlike traditional classifiers treating labels as discrete categories, vision-language models can inherently process label hierarchical structure, generalizing to unseen species through representations of related genera or families [14]. This enables robust zero-shot and few-shot learning of GeoTreeCLIP, which are critical for addressing the vast scale of global biodiversity, ever-evolving species catalogs, and the practical impossibility of exhaustive data collection for all taxa.
@@ 2.1 Open datasets for tree species classification
Table 1 provides an overview of notable open datasets for tree species classification, detailing their geographic coverage, size, taxonomic diversity, and publication year. Several datasets have contributed to tree species classification. For instance, the Seu Nico Forest dataset [15] from Brazil provides geolocated samples for 228 species but is geographically constrained. Similarly, the Maraca Ecological Station dataset [16] includes 110 species but is also region-specific. In Europe, the EUForest dataset [17] offers broader coverage with data for 242 species. On a global scale, datasets such as Tallo [18] provide significant taxonomic diversity, covering 5,163 species across 187 families. However, these datasets lack integration with remote sensing or environmental variables, limiting their application in ecological modeling.
@@
Seu Nico Forest [15] Brazil 2,868 54 families; 139 genera; 228 species 2015
EUForest [17] Europe 588,983 83 genera; 242 species 2017
Maraca Eco. Sta. [16] Brazil 680 40 families; 110 species 2020
TreeSatAI [19] Germany 50,381 15 genera; 20 species 2022
Tallo [18] Global 498,839 187 families; 1,453 genera; 5,163 species 2022
Indi. Tree Point Clouds [20] Germany 1,491 22 species 2022
NEON Veg. Struc. [21] USA N/A 949 genera; 2,826 species 2023
PureForest [22] France 135,569 18 species 2024
Planted [23] Global 2,264,747 46 genera; 40 species 2024
@@
Advances in high-resolution imaging and lidar technologies have enabled datasets like PureForest [22] and Individual Tree Point Clouds [20], which utilize aerial and point cloud data for species classification. While these datasets offer detailed structural information, they remain region-specific and lack the spectral and temporal depth of satellite-based datasets. The TreeSatAI dataset [19] combines multi-sensor data, including aerial imagery and Sentinel-1/2, for tree species classification in Germany but covers only 20 species. Similarly, the Planted dataset [23] focuses on only 40 planted species globally, limiting its broader applicability.
While valuable, these existing datasets highlight the persistent need for a benchmark that synergizes global coverage, deep taxonomic information, and multimodal remote sensing data, a gap GlobalGeoTree aims to fill.
@@ 2.2 Vision-language models for remote sensing applications
Vision-language models (VLMs) enable the integration of visual and textual information. Among these, Contrastive Language-Image Pretraining (CLIP) [13] has demonstrated exceptional zero-shot transfer capabilities by jointly training image and text encoders through a contrastive learning objective, aligning image-text pairs within a shared embedding space. For tree species classification, CLIP’s ability to learn from image-text pairings (such as satellite imagery and taxonomic labels) offers a path to capture complex visual and semantic relationships. Its proven zero-shot capabilities are particularly suited for addressing the challenges of identifying species within dynamic and evolving catalogs [14]. Meanwhile, its few-shot capabilities tackle the issue of limited labeled data, a common obstacle in biodiversity research, offering an advantage over traditional supervised methods.
In remote sensing, VLMs have been applied to various tasks, including image classification, retrieval, and scene understanding, with domain-specific adaptations yielding significant improvements. For example, RemoteCLIP [24], the first VLM specifically tailored for remote sensing, leverages pretraining on large-scale remote sensing imagery paired with aligned text, achieving state-of-the-art performance in zero-shot classification, linear probing, and few-shot learning. Similarly, SkyCLIP [25] and GeoLangBind [26] extend the capabilities of CLIP through continual pretraining on semantically diverse remote sensing image-text pairs. These models demonstrate enhanced generalization and transferability, achieving substantial gains in tasks such as zero-shot scene classification, fine-grained classification, and cross-modal retrieval compared to the original CLIP model.
@@ 3.1 Geolocated data collection and preprocessing
Tree species catalog construction
@@
We constructed a comprehensive tree species catalog by integrating two major global repositories: TreeGOER [27] and GlobalTreeSearch [28], containing 48,129 and 57,681 tree species respectively. This compilation was further enriched with multiple open-source datasets documented in Table 1. The taxonomic framework was subsequently validated and expanded using the Global Biodiversity Information Facility (GBIF) Species API [29], ensuring nomenclatural consistency and accuracy. The resulting catalog encompasses 87,845 species, representing the global diversity of tree species.
Geolocation sampling
@@
For each tree species in our catalog, we queried the GBIF Occurrence API [30] to retrieve global geolocations with documented occurrences. To ensure data quality and reliability, we applied strict filtering criteria, including: (1) selecting only recent observations recorded between 2015 and 2024; (2) limiting data to human observation records; (3) excluding records with geospatial issues as flagged by GBIF (e.g., country-coordinate mismatches); (4) filtering for occurrences with a "present" status; and (5) removing duplicate entries and observations with low geographic precision. Additionally, we ensured all samples conform to open data licenses (CC0 1.0, CC-BY-4.0, etc.), maintaining the dataset’s accessibility and reusability for the broader research community.
Forest layer filtering
To ensure that our dataset focuses on actual forest areas rather than street trees, parks, or urban vegetation, we utilized the EC JRC Global Map of Forest Cover 2020 [31] at 10m resolution. This filtering step eliminated potential samples from non-forest environments, improving the dataset’s relevance for forestry and ecological applications.
@@ 3.2 Paired remote sensing data
The resulting GlobalGeoTree dataset comprises 6,263,345 samples distributed across 221 countries and regions. More details can be checked in Appendix A. Table 2 provides an overview of the features included in each sample, which consist of paired Earth Observation (EO) data and auxiliary environmental variables derived from remote sensing sources.
@@
EO data
The Sentinel-2 data for each sample consists of a time series of 12 monthly median composites from January to December 2020. For each month, all L2A Sentinel-2 images with less than 30% cloud cover were collected, and the median composites were generated from these images. For each composite, a 5×5 pixel patch centered on the geolocation of the tree species is included. The selected patch size accounts for typical crown sizes (10–30 m) [18] and aligns with other public datasets (e.g., PureForest [22], TreeSatAI [19]), providing valuable spatial context around the specific location. The full-year temporal coverage enables models to capture the phenological patterns exhibited by different tree species across seasons [5].
Auxiliary data
The auxiliary data enriches the EO data for each sample by providing additional contextual environmental information. Geographic variables, such as elevation, slope, and aspect, are derived from the USGS SRTM dataset [32], while soil data, including volumetric water content at 33 kPa across three depths (0–5 cm, 5–15 cm, 15–30 cm), are obtained from SoilGrid [33]. Additionally, 19 bioclimatic variables are sourced from WorldClim [34]. Due to the relatively coarse spatial resolution of these datasets (ranging from 30m to 1km), only the values corresponding to the exact coordinates of each occurrence are extracted to ensure precision and relevance [35].
@@ 3.3 Dataset partitioning
Figure 3: Species in GlobalGeoTree are categorized into Frequent, Common and Rare groups based on the samples per species.
GlobalGeoTree-10kEval, is a carefully curated dataset intended for benchmarking model performance across taxonomic levels and species frequency categories in a fair and robust manner. To address the characteristic long-tail distribution observed in the datasets (detailed in Appendix A.2), species in GlobalGeoTree dataset were categorized into three groups based on the number of samples: Frequent (more than 1500), Common (100–1500), and Rare (less than 100), as shown in Fig. 3.
@@
The GlobalGeoTree-10kEval dataset includes 30 species from each of these three categories, resulting in a total of 90 species. The sample proportions within this evaluation set are 12% for Rare species, 33% for Common species, and 55% for Frequent species, culminating in around 10,000 samples. Fig. 2 shows the geographical distribution of GlobalGeoTree-10kEval, which spans diverse regions across the globe. This global distribution ensures that the dataset captures a wide range of ecological and environmental contexts, making it representative of real-world scenarios. By focusing on a diverse set of species with varying levels of representation, GlobalGeoTree-10kEval serves as a robust evaluation benchmark for assessing the ability of models to tackle challenges posed by the long-tail distribution of species and shifts in geographical domains.
@@
Beyond GlobalGeoTree-10kEval, we developed larger evaluation sets—GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900—by selecting 100 and 300 species from each category, respectively. Crucially, all samples within these evaluation sets, are excluded from the GlobalGeoTree-6M pretraining set to ensure fair evaluation. Details of all evaluation subsets (Appendix B) and the corresponding evaluation results (Appendix C) are also provided. Given the complexity of global tree species classification, our primary analysis focuses on the 90-species GlobalGeoTree-10kEval, which serves as a practical starting point for systematic benchmarking.

### v3

@@ Abstract
Global tree species mapping using remote sensing data is vital for biodiversity monitoring, forest management, and ecological research. However, progress in this field has been constrained by the scarcity of large-scale, labeled datasets. To address this, we introduce GlobalGeoTree, a comprehensive global dataset for tree species classification. GlobalGeoTree comprises 6.3 million geolocated tree occurrences, spanning 275 families, 2,734 genera, and 21,001 species across the hierarchical taxonomic levels. Each sample is paired with Sentinel-2 image time series and 27 auxiliary environmental variables, encompassing bioclimatic, geographic, and soil data. The dataset is partitioned into GlobalGeoTree-6M, a large subset for model pretraining, and curated evaluation subsets, primarily GlobalGeoTree-10kEval, a benchmark for zero-shot and few-shot classification. To demonstrate the utility of the dataset, we introduce a baseline model, GeoTreeCLIP, which leverages paired remote sensing data and taxonomic text labels within a vision-language framework pretrained on GlobalGeoTree-6M. Experimental results show that GeoTreeCLIP achieves substantial improvements in zero- and few-shot classification on GlobalGeoTree-10kEval over existing advanced models. By making the dataset, models, and code publicly available, we aim to establish a benchmark to advance tree species classification and foster innovation in biodiversity research and ecological applications. The code is publicly available at https://github.com/MUYang99/GlobalGeoTree, and the GlobalGeoTree dataset is available at https://huggingface.co/datasets/yann111/GlobalGeoTree (Mu et al., 2025b).
1 Introduction
Forests cover approximately 31% of the global land surface (Hansen et al., 2013) and provide essential ecosystem services, including carbon sequestration (Jenkins et al., 2003), biodiversity conservation (Lindenmayer et al., 2006), and climate regulation (Bonan, 2008). Accurate and large-scale mapping of tree species plays an increasingly vital role in addressing pressing environmental challenges (Mu et al., 2025a), including effective biodiversity monitoring (Felton et al., 2020), informed forest management practices (Franklin, 2001), and comprehensive ecological research aimed at understanding the complex impacts of climate change (Hamann and Wang, 2006).
@@ 1 Introduction
Traditional ground-based forest monitoring methods (Wellbrock et al., 2018), while providing detailed information, are often limited in their spatial and temporal coverage, making it challenging to obtain a comprehensive understanding of global forest composition and dynamics. In contrast, remote sensing has emerged as a key technology for large-scale forest monitoring, offering non-invasive and cost-effective approaches to tree species classification (Hermosilla et al., 2022). Despite significant advancements in this field, progress has been constrained by the limited availability of comprehensive, high-quality, and accurately labeled datasets that capture the global diversity of tree species (Bountos et al., 2025). Existing datasets typically focus on specific geographic regions or limited taxonomic coverage, hampering the development of models with global applicability (Ouaknine et al., 2025).
@@
In addition to the dataset, we introduce GeoTreeCLIP, a vision-language model specifically designed for tree species classification. Drawing on frameworks like CLIP (Radford et al., 2021), our approach aligns satellite imagery with taxonomic labels to learn nuanced representations. Unlike traditional classifiers treating labels as discrete categories, vision-language models can inherently process label hierarchical structure, generalizing to unseen species through representations of related genera or families (Stevens et al., 2024). This enables robust zero-shot and few-shot learning of GeoTreeCLIP, which are critical for addressing the vast scale of global biodiversity, ever-evolving species catalogs, and the practical impossibility of exhaustive data collection for all taxa.
@@ 2.1 Open datasets for tree species classification
Table 1 provides an overview of notable open datasets that have contributed to tree species classification, detailing their geographic coverage, size, and taxonomic diversity. For instance, the Seu Nico Forest dataset (Gastauer et al., 2015) from Brazil provides geolocated samples for 228 species but is geographically constrained. Similarly, the Maraca Ecological Station dataset (Farias et al., 2020) includes 110 species but is also region-specific. In Europe, the EUForest dataset (Mauri, 2017) offers broader coverage with data for 242 species. On a global scale, datasets such as Tallo (Jucker et al., 2022) provide significant taxonomic diversity, covering 5,163 species across 187 families. However, these datasets lack integration with remote sensing or environmental variables, limiting their application in ecological modeling.
@@
Seu Nico Forest (Gastauer et al., 2015) Brazil 2,868 54 families; 139 genera; 228 species 2015
EUForest (Mauri, 2017) Europe 588,983 83 genera; 242 species 2017
Maraca Eco. Sta. (Farias et al., 2020) Brazil 680 40 families; 110 species 2020
TreeSatAI (Ahlswede, 2022) Germany 50,381 15 genera; 20 species 2022
Tallo (Jucker et al., 2022) Global 498,839 187 families; 1,453 genera; 5,163 species 2022
Indi. Tree Point Clouds (Weiser et al., 2022) Germany 1,491 22 species 2022
NEON Veg. Struc. (Kampe et al., 2010) USA N/A 949 genera; 2,826 species 2023
PureForest (Gaydon and Roche, 2025) France 135,569 18 species 2024
Planted (Pazos-Outón et al., 2024) Global 2,264,747 46 genera; 40 species 2024
@@
Advances in high-resolution imaging and lidar technologies have enabled datasets like PureForest (Gaydon and Roche, 2025) and Individual Tree Point Clouds (Weiser et al., 2022), which utilize aerial and point cloud data for species classification. While these datasets offer detailed structural information, they remain region-specific and lack the spectral and temporal depth of satellite-based datasets. The TreeSatAI dataset (Ahlswede, 2022) combines multi-sensor data, including aerial imagery and Sentinel-1/2, for tree species classification in Germany but covers only 20 species. Similarly, the Planted dataset (Pazos-Outón et al., 2024) focuses on only 40 planted species globally, limiting its broader applicability. Our work also complements broader biodiversity benchmarks like GeoLifeCLEF (Botella et al., 2025) and GeoPlant (Picek et al., 2024). While these focus on regional, multi-lifeform plant prediction, GlobalGeoTree is the first dataset with a global scope specifically curated for over 21,000 forest tree species.
Collectively, while valuable, these existing datasets highlight the persistent need for a benchmark that synergizes global coverage, deep taxonomic information for forest tree species, and multimodal remote sensing data, a gap GlobalGeoTree aims to fill.
@@ 2.2 Vision-language models for remote sensing applications
Vision-language models (VLMs) enable the integration of visual and textual information. Among these, Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) has demonstrated exceptional zero-shot transfer capabilities by jointly training image and text encoders through a contrastive learning objective, aligning image-text pairs within a shared embedding space. For tree species classification, CLIP’s ability to learn from image-text pairings (such as satellite imagery and taxonomic labels) offers a path to capture complex visual and semantic relationships. Its proven zero-shot capabilities are particularly suited for addressing the challenges of identifying species within dynamic and evolving catalogs (Stevens et al., 2024). Meanwhile, its few-shot capabilities tackle the issue of limited labeled data, a common obstacle in biodiversity research, offering an advantage over traditional supervised methods.
In remote sensing, VLMs have been applied to various tasks, including image classification, retrieval, and scene understanding, with domain-specific adaptations yielding significant improvements. For example, RemoteCLIP (Liu et al., 2024), the first VLM specifically tailored for remote sensing, leverages pretraining on large-scale remote sensing imagery paired with aligned text, achieving state-of-the-art performance in zero-shot classification, linear probing, and few-shot learning. Similarly, SkyCLIP (Wang et al., 2024), SkyCLIP-50 (Wang et al., 2024) and CLIP-laion-RS (He et al., 2024) extend the capabilities of CLIP through continual pretraining on semantically diverse remote sensing image-text pairs. These models demonstrate enhanced generalization and transferability, achieving substantial gains in tasks such as zero-shot scene classification, fine-grained classification, and cross-modal retrieval compared to the original CLIP model.
@@ 3.1 Geolocated data collection and preprocessing
3.1.1 Tree species catalog construction
@@
We constructed a comprehensive tree species catalog by integrating two major global repositories: TreeGOER (Kindt, 2023) and GlobalTreeSearch (Beech et al., 2017), containing 48,129 and 57,681 tree species respectively. This compilation was further enriched with multiple open-source datasets documented in Table 1. The taxonomic framework was subsequently validated and expanded using the Global Biodiversity Information Facility (GBIF) Species API (Global Biodiversity Information Facility (2025), GBIF), ensuring nomenclatural consistency and accuracy. The resulting catalog encompasses 87,845 species, representing the global diversity of tree species.
3.1.2 Geolocation sampling
@@
For each tree species in our catalog, we queried the GBIF Occurrence API (Chamberlain and Boettiger, 2017) to retrieve global geolocations with documented occurrences. To ensure data quality and reliability, we applied strict filtering criteria, including: (1) selecting only recent observations recorded between 2015 and 2024; (2) limiting data to human observation records; (3) excluding records with geospatial issues as flagged by GBIF (e.g., country-coordinate mismatches); (4) filtering for occurrences with a "present" status; and (5) removing duplicate entries and observations with low geographic precision. Additionally, we ensured all samples conform to open data licenses (CC0 1.0, CC-BY-4.0, etc.), maintaining the dataset’s accessibility and reusability for the broader research community.
3.1.3 Forest layer filtering
To ensure that each geolocated observation corresponds to a valid tree, we performed an additional forest cover verification step. We utilized the EC JRC Global Map of Forest Cover 2020 (Version 2) (Bourgoin et al., 2025), which has a 10m spatial resolution. Each geolocated point from GBIF was cross-referenced with this map, and only samples located within forest areas were retained. This filtering not only served as an essential quality control measure to enhance data reliability but also defined 2020 as the target year for our study. This allowed us to subsequently acquire the Sentinel-2 time series for 2020, ensuring a direct temporal correspondence between the verified ground observations and the satellite imagery. While we utilized the binary JRC mask for this study, future operational pipelines could integrate multi-class land cover products such as GLC_FCS10 (Zhang et al., 2025) to constrain candidate species based on mapped forest types.
3.2 Dataset Composition and Structure
The filtering process results in the final GlobalGeoTree dataset, which comprises 6,263,345 high-quality samples distributed across 221 countries and regions. Approximately 70% of the records are sourced from "Research Grade" iNaturalist observations, which require identification consensus from at least two independent contributors. Other samples are mainly from authoritative datasets available through GBIF, including National Forest Inventory (NFI) data from countries such as Sweden, France, and Colombia. The full list of source datasets and their provenance is traceable via the unique GBIF Derived Dataset DOI of GlobalGeoTree: https://doi.org/10.15468/dd.9qxqyy. A core feature of this dataset is its multi-granular taxonomic hierarchy, which is essential for developing models that can generalize across different levels of biological classification. This hierarchical structure is visualized in Fig. 2, illustrating the relationships branching from four broad functional types at the center (level 0), through 275 families (level 1) and 2,734 genera (level 2), out to the 21,001 individual species at the periphery (level 3).
Figure 2: The taxonomic hierarchy of the GlobalGeoTree dataset. The visualization shows the nested relationships, branching from the four functional types (e.g., Deciduous Broadleaf) at the center, through families (e.g., Fabaceae), genera, and out to the 21,001 species at the outermost ring. This multi-level structure is a core feature that enables multi-granular classification tasks.
Each of the 6.3 million samples is structured as a single record that includes this full taxonomic lineage, its geolocation, and source metadata. This entire collection is made available as a downloadable CSV file (GlobalGeoTree.csv) for full transparency and accessibility. Table 2 presents several example records to illustrate the data structure. A comprehensive statistical breakdown and accessibility details of the dataset is provided in Appendix A.
Table 2: Example records from the GlobalGeoTree.csv file, demonstrating the structure of each geolocated data point.
country_code level0 level1_family level2_genus level3_species source longitude latitude
ES Deciduous Broadleaf Fagaceae Quercus Quercus rotundifolia iNaturalist Res.-grade Obs. -3.74024 40.49546
US Evergreen Needleleaf Cupressaceae Thuja Thuja plicata iNaturalist Res.-grade Obs. -122.168 48.33392
IE Deciduous Broadleaf Fagaceae Castanea Castanea sativa iNaturalist Res.-grade Obs. -9.524 52.05556
PT Deciduous Broadleaf Fagaceae Quercus Quercus rotundifolia Forestry Inventory 2015 -7.85145 39.42701
AU Deciduous Broadleaf Fabaceae Acacia Acacia platycarpa ALA species sightings 132.9895 -14.818
@@ 3.3 Paired remote sensing data
Each sample in the GlobalGeoTree dataset is paired with a rich set of multimodal data to enable robust modeling. Table 3 provides a comprehensive overview of these features, including paired Earth Observation (EO) data and auxiliary environmental variables derived from remote sensing sources.
@@
3.3.1 EO data
The Sentinel-2 data for each sample consists of a time series of 12 monthly median composites from January to December 2020. The full-year temporal coverage enables models to capture the distinct seasonal patterns of different tree species, such as leaf-out, full canopy, and senescence (Mu et al., 2025a). These phenological features are particularly discriminative for deciduous species.
For each month, all L2A Sentinel-2 images with less than 30% cloud cover were collected, and the median composites were generated from these images. This process ensures a robust, largely cloud-free representation for each month, effectively mitigating issues caused by transient atmospheric conditions in single-date imagery.
Each composite includes a 5×5 pixel patch centered on the geolocation of the tree species. This patch size was chosen to account for typical crown sizes (10–30 m) (Jucker et al., 2022) and to align with other public datasets, such as PureForest (Gaydon and Roche, 2025) and TreeSatAI (Ahlswede, 2022). The chosen 5×5 patch size represents a trade-off: it is sufficiently large enough to capture the crown of a mature tree and its immediate context, while remaining compact enough to minimize noise from surrounding unrelated species.
3.3.2 Auxiliary data
The auxiliary data enriches the EO data for each sample by providing additional contextual environmental information. Geographic variables, such as elevation, slope, and aspect, are derived from the USGS SRTM dataset (Jarvis et al., 2008), while soil data, including volumetric water content at 33 kPa across three depths (0–5 cm, 5–15 cm, 15–30 cm), are obtained from SoilGrid (Poggio et al., 2021). Additionally, 19 bioclimatic variables are sourced from WorldClim (Fick and Hijmans, 2017). Due to the relatively coarse spatial resolution of these datasets (ranging from 30m to 1km), only the values corresponding to the exact coordinates of each occurrence are extracted to ensure precision and relevance (Gillespie et al., 2024). A detailed description of all 27 auxiliary variables and their sources is provided in Appendix B.
@@ 3.4 Dataset partitioning
Figure 3: Species in GlobalGeoTree are categorized into Frequent, Common and Rare groups based on the number of samples per species.
GlobalGeoTree-10kEval is a carefully curated dataset designed to benchmark model performance across taxonomic levels and species frequency categories. To address the characteristic long-tail distribution (detailed in Appendix A.3), we categorized species into three groups based on sample frequency: Frequent (>1500>1500 samples), Common ($100\text{--}1500$ samples), and Rare (<100<100 samples), as shown in Fig. 3.
@@
The primary GlobalGeoTree-10kEval dataset includes 30 species from each of these three categories, resulting in a total of 90 species. The sample proportions within this evaluation set are 12% for Rare species, 33% for Common species, and 55% for Frequent species, culminating in around 10,000 samples. Fig. 4 shows the geographical distribution of GlobalGeoTree-10kEval, which spans diverse regions across the globe. This global distribution ensures that the dataset captures a wide range of ecological and environmental contexts, making it representative of real-world scenarios. By focusing on a diverse set of species with varying levels of representation, GlobalGeoTree-10kEval serves as a robust evaluation benchmark for assessing the ability of models to tackle challenges posed by the long-tail distribution of species and shifts in geographical domains.
@@
To further evaluate model robustness and scalability across broader taxonomic scopes, we constructed two additional evaluation subsets: GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900, containing 100 and 300 species per category, respectively. Crucially, all samples within these evaluation sets, were subjected to the rigorous filtering criteria detailed in Sect. 3.1.2 and are strictly excluded from the GlobalGeoTree-6M pretraining set to ensure fair evaluation. Details of all evaluation subsets (Appendix C) and the corresponding evaluation results (Appendix D) are also provided. Given the complexity of global tree species classification, our primary analysis focuses on the 90-species GlobalGeoTree-10kEval, which serves as a practical starting point for systematic benchmarking.
4 Data Quality and Validation
Figure 5: The GlobalGeoTree Explorer App. The interface allows users to visualize the global distribution of tree samples (bottom map) and inspect individual points against high-resolution satellite imagery (top panels a-f). Detailed attributes for selected samples are displayed in the sidebar, facilitating transparent community validation.
To ensure the reliability of GlobalGeoTree, we implemented a multi-tiered validation framework combining fine-scale visual inspection, global-scale cross-verification with independent land cover products, and a public platform for community auditing.
4.1 Visual Inspection and Community Validation
We first quantified land-cover accuracy through a visual validation study. A random subset of 300 locations was sampled from GlobalGeoTree-10kEval. For each sample, we retrieved Very High Resolution (VHR) satellite imagery to verify whether the geolocated point fell within forest or tree cover. The inspection confirmed that 98.3% of the samples were correctly located in forested areas, while a small minority were ambiguous (e.g., forest edges), validating the efficacy of our automated filtering pipeline.
To further enhance transparency and enable large-scale qualitative validation by the research community, we developed the GlobalGeoTree Explorer, a web-based application hosted on Google Earth Engine (Fig. 5). This tool allows users to visualize the spatial distribution of all 6.3 million samples, overlay them on high-resolution satellite basemaps, and inspect detailed taxonomic attributes for any individual point. The application is publicly accessible at https://ee-yangm.projects.earthengine.app/view/globalgeotree-explorer.
4.2 Cross-Validation with Global Land Cover Products
To validate taxonomic consistency at the functional type level, we performed a cross-comparison against the Copernicus Global Land Cover Layers (CGLS-LC100, Collection 3) (Buchhorn et al., 2020). Since no global dataset currently provides high-resolution distribution maps at the species or genus level, the CGLS-LC100 product represents the most detailed global baseline available, offering discrete land cover layers at 100m resolution that distinguish between broadleaf and needleleaf forest types.
Figure 6: Cross-validation results against the CGLS-LC100 product. Left: Normalized confusion matrix at the Leaf Level (Needleleaf vs. Broadleaf), showing an overall agreement of 80.38%. Right: Agreement rates broken down by continent, highlighting strong consistency in tropical regions.
We aligned the level0 labels of GlobalGeoTree with CGLS classes by aggregating them into two primary leaf types: Needleleaf and Broadleaf. To ensure global representativeness and mitigate regional data density biases, we employed a stratified sampling strategy, selecting 10,000 samples per continent.
Figure 6 presents the results of this cross-validation. The confusion matrix (left) reveals an overall agreement of 80.38% at the leaf level. We observed high consistency for Broadleaf samples (81%), while some Needleleaf samples (22%) showed discrepancies with the coarser 100m product, likely due to mixed pixels in transition zones. The regional breakdown (right) demonstrates exceptionally high agreement in South America (99.3%) and Africa (98.7%), where broadleaf forests dominate. Lower agreement rates in North America and Europe are expected given the higher prevalence of mixed forests and the resolution gap between our point-based data and the 100m validation raster.
4.3 Taxonomic Reliability
Beyond spatial validation, taxonomic reliability is ensured through our strict data provenance pipeline. We exclusively utilize observations which require community consensus for species identification, and 100% of these samples have passed the GBIF taxonomic backbone check (GBIF Secretariat, 2023).